7x + 4 (x-1)=3x + 4
Answer by swincher4391(1107)

7x+ 4(x-1) = 3x+4
7x + 4x -4 = 3x+4 by distributive property
11x -4 = 3x +4
8x -4 =4
8x = 8
x =1